OpenStudy (anonymous):

Find the zeros of the function y=4x^2-11x+6 Please help!

OpenStudy (saifoo.khan):

that's same.

OpenStudy (lgbasallote):

zeros is factor then equate to zero right @saifoo.khan ?

OpenStudy (anonymous):

What're the steps to finding the zeros of this function?

OpenStudy (anonymous):

you can factorize or use the quadratic formula

OpenStudy (anonymous):

4x^2-11x+6= 4x^2-3x-8x+6=x(4x-3)-2(4x-3)=(4x-3)(x-2) so, y =0, when either 4x-3=0 ie x=3/4 or x-2=0 ie x=2

OpenStudy (anonymous):

Let \[4x ^{2}-11x+6=0\] \[4x ^{2}-8x-3x+6=0\] \[4x(x-2)-3(x-2)=0\] \[(x-2)(4x-3)=0\] Therefore either x-2=0 or 4x-3=0 x=2 or x=3/4